Consider the utility function u1(x1, x2, x3) = x21,x32 x3 .

Find a monotone transformation of u1(x1, x2, x3) such that the new utility function is of the form u3(x1, x2, x3) = a In x1 + b In x2 + c In x3.

1
Expert's answer
2021-09-29T18:05:48-0400

Such monotonic transformation can be done so that the powers are added up to 1

We see that

"2+3+1=6\\and\\a=\frac{2}{6}=\frac{1}{3}\\b=\frac{3}{6}=\frac{1}{2}\\c=\frac{1}{6}"

After monotonic transformation of "U^i(X_1,X_2,X_3)" the new utility function is "U^2(X_1,X_2,X_3)=X_1^{\frac{1}{3}},
X_2^{\frac{1}{2}}
,X_3^{\frac{1}{6}}"